a) A +100-μC charge produces 1.5 T field 2 cm from the charge. Determine the magnitude of velocity of the charge. b) What magnetic field is produced by a charge twice that in part a) 4 cm away from it?

1
Expert's answer
2022-01-26T17:46:01-0500

Part a)

"qvB=\dfrac1{4\pi\epsilon_0}\dfrac{q^2}{r^2},\\\space\\
v=\dfrac1{4\pi\epsilon_0}\dfrac{q}{r^2B}=1.5·10^9\text{ m/s}."


Part b) The field will be 4 times weaker, or 0.375 T.